# Sample Complexity of Linear Quadratic Regulator Without Initial Stability

### v2

@@ Abstract.
Inspired by REINFORCE, we introduce a novel receding-horizon algorithm for the Linear Quadratic Regulator (LQR) problem with unknown parameters. Unlike prior methods, our algorithm avoids reliance on two-point gradient estimates while maintaining the same order of sample complexity. Furthermore, it eliminates the restrictive requirement of starting with a stable initial policy, broadening its applicability. Beyond these improvements, we introduce a refined analysis of error propagation through the contraction of the Riemannian distance over the Riccati operator. This refinement leads to a better sample complexity and ensures improved convergence guarantees. Numerical simulations validate the theoretical results, demonstrating the method’s practical feasibility and performance in realistic scenarios.
@@ 1. Introduction
In contrast, the model-free approach, rooted in reinforcement learning (RL), bypasses the need for explicit dynamics identification, instead focusing on direct policy optimization through cost evaluations. Recent advances leverage stochastic zero-order optimization techniques, including policy gradient methods, to achieve provable convergence to near-optimal solutions despite the inherent non-convexity of the LQR cost landscape. Foundational works, such as [5], established the feasibility of such methods despite the non-convexity of the problem, demonstrating convergence under random initialization. Subsequent efforts, including [11] and [13], have refined these techniques, achieving improved sample complexity bounds. Notably, all of these works assume that the initial policy is stabilizing.
A key limitation of these methods, including [11, 13], is the reliance on two-point gradient estimation, which requires evaluating costs for two different policies while maintaining identical initial states. In practice, this assumption is often infeasible, as the initial state is typically chosen randomly and cannot be controlled externally. Our earlier work [12] addressed this challenge, establishing the best-known result among methods that assume initial stability without having to rely on two-point estimates. Instead, we proposed a one-point gradient estimation method, inspired by REINFORCE [19, 17], that achieves the same convergence rate as the two-point method [11] using only a single cost evaluation at each step. This approach enhances both the practical applicability and theoretical robustness of model-free methods, setting a new benchmark under the initial stability assumption.
@@
As also pointed out in [20], the $\gamma$-discounted LQR problems studied in [14, 9, 22] are equivalent to the standard non-discounted LQR with system matrices scaled by $\sqrt{\gamma}$. In [14, 9, 22], this scaling results in an enlarged set of stabilizing policies when $\gamma$ is sufficiently small, enabling policy gradient algorithms to start from an arbitrary policy. However, as noted in [20], this comes at the cost of solving multiple LQR instances rather than a single one, increasing computational overhead. Furthermore, the optimization landscape in the discounted setting remains fundamentally the same as in the undiscounted case, as described in [5, 11]. Consequently, the same difficulties mentioned in [8, 18] persist when extending these methods to output-feedback settings, where additional estimation errors complicate policy search. In contrast, receding-horizon approaches [20] provide a more direct and extensible framework for tackling such challenges [21].
This paper builds on the receding-horizon policy gradient framework introduced in [20], a significant step towards eliminating the need for a stabilizing initial policy by recursively updating finite-horizon costs. While the approach proposed in this work marks an important step forward in model-free LQR, we address the reliance on the two-point gradient estimation, a known limitation discussed earlier. Building on the gradient estimation approach from our earlier work [12], we adapt the core idea to accommodate the new setup that eliminates the initial stability assumption. Specifically, our modified method retains the same convergence rate while overcoming the restrictive assumptions of two-point estimation. Beyond these modifications, we introduce a refined convergence analysis, via an argument based on a Riemannian distance function [3], which significantly improves the propagation of errors. This ensures that the accumulated error remains linear in the horizon length, in contrast to the exponential growth in [20]. As a result, we achieve a uniform sample complexity bound of $\widetilde{\mathcal{O}}(\varepsilon^{-2})$, independent of problem-specific constants, thereby offering a more scalable and robust policy search framework.
@@ 1.1. Algorithm and Paper Structure Overview
To be more specific, the core idea of the algorithm leverages the observation that, for any error tolerance $\varepsilon$, there exists sufficiently large finite horizon $N$ where the sequence of policies minimizing recursively updated finite-horizon costs can approximate the optimal policy for the infinite-horizon cost within $\varepsilon$ neighborhood. This insight motivates the algorithm’s design: a recursive outer loop that iteratively refines the surrogate cost function over a sequence of finite horizons, and an inner loop that employs policy gradient methods to approximate the optimal policy for each of these costs. Specifically, in the outer loop, the algorithm updates the surrogate cost and the associated policy at each horizon step $h$, starting from the terminal horizon $h=N-1$ and moving backward to $h=0$. At each step $h$, the inner loop applies a policy gradient method to compute an approximately optimal policy for the finite-horizon cost over the interval $[h,N]$. This step generates a surrogate policy $\widetilde{K}_{h}$, which is then incorporated into the cost function of the subsequent step in the outer loop.
@@ 2. Preliminaries
In this section, we gather the required notation, closely following the ones in [20] which our work builds on. Consider the discrete-time linear system
@@
where $x_{t}\in{\mathbb{R}}^{n}$ is the system state at time $t$, $u_{t}\in{\mathbb{R}}^{m}$ is the control input at time $t\geq 0$, $A\in{\mathbb{R}}^{n\times n}$ and $B\in{\mathbb{R}}^{n\times m}$ are the system matrices which are unknown to the control designer. Crucially here, the initial state $x_{0}$ is sampled randomly from a distribution ${\mathcal{D}}$ and satisfies
[…]
The objective in the LQR problem is to find the optimal controller that minimizes the following cost
$J_{\infty}={\mathbb{E}}_{x_{0}\sim{\mathcal{D}}}\left[\sum_{t=0}^{\infty}x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t}\right],$
where $Q\in{\mathbb{R}}^{n\times n}$ and $R\in{\mathbb{R}}^{m\times m}$ are the symmetric positive definite matrices that parameterize the cost. We require the pair $(A,B)$ to be stabilizable, and since $Q\succ 0$, the pair $(A,Q^{1/2})$ is observable, As a result, the unique optimal controller is a linear state-feedback $u^{\ast}_{t}=-K^{*}x_{t}$ where $K^{*}$ is derived as follows
@@
and $P^{*}$ denotes the unique positive definite solution to the discounted discrete-time algebraic Riccati equation (ARE) [2]:
@@
2.1. Notations
We use $\|X\|$, $\|X\|_{F}$, $\sigma_{\min}(X)$, and $\kappa_{X}$ to denote the 2-norm, Frobenius norm, minimum singular value, and the condition number of a matrix $X$ respectively. We also use $\rho(X)$ to denote the spectral radius of a square matrix $X$. Moreover, for a positive definite matrix $W$ of appropriate dimensions, we define the $W$-induced norm of a matrix $X$ as
$\|X\|_{W}^{2}:=\sup_{z\neq 0}\frac{z^{\top}X^{\top}WXz}{z^{\top}Wz}.$
Following the notation in [20], we denote the $P^{*}$-induced norm by $\|X\|_{*}$. Furthermore, we denote the Riemannian distance [3] between two positive definite matrices $U,V\in{\mathbb{R}}^{n\times n}$ by
$\delta(U,V)=\left(\sum_{i=1}^{n}\log^{2}\lambda_{i}(UV^{-1})\right)^{1/2}.$
We now introduce some important notations which will be used in describing the algorithm and proof of the main result. Let $N$ be the horizon length and $h$ the initial time step. The true finite-horizon cost $J_{h}(K_{h})$ of a policy $K_{h}$ is defined as
(5) $J_{h}(K_{h}):={\mathbb{E}}_{x_{h}\sim\mathcal{D}}\Bigg{[}\sum_{t=h+1}^{N-1}x_{t}^{\top}\left(Q+(K^{*}_{t})^{\top}RK^{*}_{t}\right)x_{t}+x_{h}^{\top}\left(Q+K_{h}^{\top}RK_{h}\right)x_{h}+x_{N}^{\top}Q_{N}x_{N}\Bigg{]},$
@@ 2.1. Notations
  • $x_{h}\sim\mathcal{D}$ denotes the initial state $x_{h}$ is drawn from the distribution $\mathcal{D}$,
  • $Q_{N}$ is the terminal cost matrix, which can be chosen arbitrarily (e.g., $Q_{N}=0$),
@@
(6) $\widetilde{J}_{h}(K_{h}):={\mathbb{E}}_{x_{h}\sim\mathcal{D}}\Bigg{[}\sum_{t=h+1}^{N-1}x_{t}^{\top}\left(Q+\widetilde{K}_{t}^{\top}R\widetilde{K}_{t}\right)x_{t}+x_{h}^{\top}\left(Q+K_{h}^{\top}RK_{h}\right)x_{h}+x_{N}^{\top}Q_{N}x_{N}\Bigg{]},$
@@
$x_{h+1}=(A-BK_{h})x_{h}.$
The key difference between $\widetilde{J}_{h}(K_{h})$ and $J_{h}(K_{h})$ lies in the use of $\widetilde{K}_{t}$ versus $K^{*}_{t}$ for $t\in\{h+1,\dotsc,N-1\}$. This distinction implies that $\widetilde{J}_{h}(K_{h})$ incorporates all errors from earlier steps, precisely the ones at $\{N-1,\dotsc,h+1\}$, as the procedure is recursive.
@@
(7) $\widetilde{J}_{h}(K_{h};x_{h})$ $:=\sum_{t=h+1}^{N-1}x_{t}^{\top}\left(Q+\widetilde{K}_{t}^{\top}R\widetilde{K}_{t}\right)x_{t}+x_{h}^{\top}(Q+K_{h}^{\top}RK_{h})x_{h}+x_{N}^{\top}Q_{N}x_{N}$
(8) $=x_{h}^{\top}(Q+K_{h}^{\top}RK_{h})x_{h}+x_{h}^{\top}(A-BK_{h})^{\top}\widetilde{P}_{h+1}(A-BK_{h})x_{h},$
@@
$\widetilde{J}_{h}(K_{h})={\mathbb{E}}_{x_{h}\sim\mathcal{D}}\left[\widetilde{J}_{h}(K_{h};x_{h})\right].$
Using (8), we can compute the gradient of $\widetilde{J}_{h}(K_{h};x_{h})$ with respect to $K_{h}$ as follows:
(9) $\nabla\widetilde{J}_{h}(K_{h};x_{h})$ $=\nabla\Bigg{(}x_{h}^{\top}K_{h}^{\top}RK_{h}x_{h}+x_{h}^{\top}K_{h}^{\top}B^{\top}\widetilde{P}_{h+1}BK_{h}x_{h}-2x_{h}^{\top}A^{\top}\widetilde{P}_{h+1}BK_{h}x_{h}\Bigg{)}$
(10) $=2RK_{h}x_{h}x_{h}^{\top}+2B^{\top}\widetilde{P}_{h+1}BK_{h}x_{h}x_{h}^{\top}-2B^{\top}\widetilde{P}_{h+1}Ax_{h}x_{h}^{\top}$
(11) $=2\left((R+B^{\top}\widetilde{P}_{h+1}B)K_{h}-B^{\top}\widetilde{P}_{h+1}A\right)x_{h}x_{h}^{\top},$
@@
(12) $\nabla\widetilde{J}_{h}(K_{h})$ $={\mathbb{E}}_{x_{h}\sim\mathcal{D}}\left[\nabla\widetilde{J}_{h}(K_{h};x_{h})\right]$
(13) $=2\left((R+B^{\top}\widetilde{P}_{h+1}B)K_{h}-B^{\top}\widetilde{P}_{h+1}A\right){\mathbb{E}}_{x_{h}\sim\mathcal{D}}\left[x_{h}x_{h}^{\top}\right]$
@@
$Q_{h}(x_{h},u_{h})$ $:=x_{h}^{\top}Qx_{h}+u_{h}^{\top}Ru_{h}+\sum_{t=h+1}^{N-1}x_{t}^{\top}\left(Q+\widetilde{K}_{t}^{\top}R\widetilde{K}_{t}\right)x_{t}+x_{N}^{\top}Q_{N}x_{N}$
$=x_{h}^{\top}Qx_{h}+u_{h}^{\top}Ru_{h}+\widetilde{J}_{h+1}(\widetilde{K}_{h+1};Ax_{h}+Bu_{h})$
(15) $=x_{h}^{\top}Qx_{h}+u_{h}^{\top}Ru_{h}+(Ax_{h}+Bu_{h})^{\top}\widetilde{P}_{h+1}(Ax_{h}+Bu_{h}),$
@@
$\widetilde{J}_{h}(K_{h})={\mathbb{E}}_{x_{h}\sim\mathcal{D}}\left[Q_{h}(x_{h},-K_{h}x_{h})\right].$
@@
(22) ${\mathbb{E}}_{x_{h}\sim{\mathcal{D}}}\left[x_{h}^{\top}\widetilde{P}_{h}x_{h}\right]$ $=\widetilde{J}_{h}(\widetilde{K}_{h}),$
(23) ${\mathbb{E}}_{x_{h}\sim{\mathcal{D}}}\left[x_{h}^{\top}\widetilde{P}^{*}_{h}x_{h}\right]$ $=\widetilde{J}_{h}(\widetilde{K}^{*}_{h}),\quad\text{and}$
(24) ${\mathbb{E}}_{x_{h}\sim{\mathcal{D}}}\left[x_{h}^{\top}P^{*}_{h}x_{h}\right]$ $=J_{h}(K^{*}_{h}).$
@@
We now introduce the following mild assumption, which will be useful in establishing a key result.
@@
$A$ in (1) is non-singular.
Under this assumption, the following result from [16] holds:

### v3

@@ Abstract.
Inspired by REINFORCE, we introduce a novel receding-horizon algorithm for the Linear Quadratic Regulator (LQR) problem with unknown dynamics. Unlike prior methods, our algorithm avoids reliance on two-point gradient estimates while maintaining the same order of sample complexity. Furthermore, it eliminates the restrictive requirement of starting with a stable initial policy, broadening its applicability. Beyond these improvements, we introduce a refined analysis of error propagation through the contraction of the Riccati operator under the Riemannian distance. This refinement leads to a better sample complexity and ensures improved convergence guarantees.
@@ 1. Introduction
In contrast, the model-free approach, rooted in reinforcement learning (RL), bypasses the need for explicit dynamics identification, instead focusing on direct policy optimization through cost evaluations. Recent advances leverage stochastic zero-order optimization techniques, including policy gradient methods, to achieve provable convergence to near-optimal solutions despite the inherent non-convexity of the LQR cost landscape. Foundational works, such as [6], established the feasibility of such methods, demonstrating convergence to the globally optimal solution. Subsequent efforts, including [14] and [16], have refined these techniques, achieving improved sample complexity bounds. Notably, all of these works assume that the initial policy is stabilizing.
A key limitation of these methods, including [14, 16], is the reliance on two-point gradient estimation, which requires evaluating costs for two different policies while maintaining identical initial states. In practice, this assumption is often infeasible, as the initial state is typically chosen randomly and cannot be controlled externally. Our earlier work [15] addressed this challenge, establishing the best-known result among methods that assume initial stability without having to rely on two-point estimates. Instead, we proposed a gradient estimation method inspired by REINFORCE [21, 20], that achieves the same convergence rate as the two-point method [14] while using only a single cost evaluation at each step. This approach enhances the practical applicability of model-free methods, setting a new benchmark under the initial stability assumption.
@@
As also pointed out in [22], the $\gamma$-discounted LQR problems studied in [17, 12, 23] are equivalent to the standard undiscounted LQR with system matrices scaled by $\sqrt{\gamma}$. In [17, 12, 23], this scaling results in an enlarged set of stabilizing policies when $\gamma$ is sufficiently small, enabling policy gradient algorithms to start from an arbitrary policy by choosing $\gamma$ appropriately. However, in this setting, the discounting factor serves as an additional parameter that can be tuned (e.g., as in [17]). In contrast, receding-horizon approaches [22], and our work here, consider the undiscounted LQR setting, where no discount factor is introduced, and if one were present, it would be treated as a fixed problem parameter rather than a tunable one.
This paper builds on the receding-horizon policy gradient framework introduced in [22], a significant step towards eliminating the need for a stabilizing initial policy by recursively updating finite-horizon costs. While the approach proposed in [22] marks an important step forward in model-free LQR, we address the reliance on the two-point gradient estimation, a known limitation discussed earlier. Building on the gradient estimation approach from our earlier work [15], we adapt the core idea to accommodate the new setup that eliminates the initial stability assumption. Specifically, our modified method retains the same convergence rate while overcoming the restrictive assumptions of two-point estimation. Beyond these modifications, we revisit the convergence analysis in [22], where a calculation mistake obscured the fact that their error accumulation grows exponentially with the horizon length, invalidating their claimed convergence guarantee. Our refined argument, based on a Riemannian distance function [3], significantly improves the propagation of errors. This ensures that the accumulated error remains linear in the horizon length, in contrast to the exponential growth in [22]. As a result, we achieve a uniform sample complexity bound of $\widetilde{\mathcal{O}}(\varepsilon^{-2})$, independent of problem-specific constants in the exponent, thereby offering a more scalable and robust policy search framework.
In summary, the key contributions of this work are:
  • In addition to correcting the required sample complexity of [22], we develop a new outer loop analysis (under a mild additional assumption) that establishes convergence guarantees for much smaller sample complexity rates than the actual rate in [22].
  • We propose a more practical gradient estimate that avoids reliance on two-point estimation and its assumption of the ability to select the randomness of system.
@@ 1.1. Algorithm and Paper Structure Overview
To be more specific, the core idea of the algorithm leverages the observation that, for any error tolerance $\varepsilon$, there exists sufficiently large finite horizon $N$ where the sequence of policies minimizing recursively updated finite-horizon costs can approximate the optimal policy for the infinite-horizon cost within $\varepsilon$ neighborhood. This receding-horizon design is key to eliminating the need for an initial stabilizing policy: because each horizon step solves a finite-horizon problem, all stage costs remain finite even when the initial policy is unstable. This insight motivates the algorithm’s design: a backward outer loop that iteratively constructs surrogate cost functions over a sequence of finite horizons, and a forward inner loop that optimizes these surrogate costs. At each stage $h$, the policies for later steps $h+1$ to N1N-1 are frozen, and the inner loop applies a policy gradient method to optimize only over the policy at stage $h$, yielding an approximately optimal policy $\widetilde{K}_{h}$ for the finite-horizon cost for the interval $[h,N]$. The resulting policy $\widetilde{K}_{h}$ is then incorporated into the cost function of the subsequent step in the outer loop, and the process repeats backward until $h=0$. Finally, note that the inner loop uses a model-free, one-point gradient estimator, which relies solely on observed costs and does not require knowledge of the system dynamics.
@@ 2. Preliminaries
In this section, we gather the required notation, closely following the ones in [22] which our work builds on. We use $\|X\|$, $\|X\|_{F}$, $\sigma_{\min}(X)$, and $\kappa_{X}$ to denote the 2-norm, Frobenius norm, minimum singular value, and the condition number of a matrix $X$ respectively. We also use $\rho(X)$ to denote the spectral radius of a square matrix $X$. Moreover, for a positive definite matrix $W$ of appropriate dimensions, we define the $W$-induced norm of a matrix $X$ as
$\|X\|_{W}^{2}:=\sup_{z\neq 0}\frac{z^{\top}X^{\top}WXz}{z^{\top}Wz}.$
Furthermore, we denote the Riemannian distance [3] between two positive definite matrices $U,V\in{\mathbb{R}}^{n\times n}$ by
$\delta(U,V):=\left(\sum_{i=1}^{n}\log^{2}\lambda_{i}(UV^{-1})\right)^{1/2}.$
Now consider the discrete-time linear system
@@
where $x_{t}\in^{n}$ is the system state at time $t\in\{0,1,2,\dotsc\}$, $u_{t}\in^{m}$ is the control input at time $t$, $A\in^{n\times n}$ and $B\in^{n\times m}$ are the system matrices which are unknown to the control designer. Crucially here, the initial state $x_{0}$ is drawn randomly from a distribution $\mathcal{D}$} and satisfies
(2) ${\mathbb{E}}[x_{0}]=0,\quad{\mathbb{E}}[x_{0}x_{0}^{\top}]=\Sigma_{0},\ \text{and}\ \|x_{0}\|^{2}\leq C_{m}\ \text{a.s.}$
for some positive definite $\Sigma_{0}\in^{n\times n}$ and scalar $C_{m}>0$. This assumption of a random initial state has been widely adopted in the recent LQR literature [6, 14, 16, 10, 8, 13, 4, 11]. The objective in the LQR problem is to find the optimal controller that minimizes the following cost
$J_{\infty}={\mathbb{E}}\left[\sum_{t=0}^{\infty}x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t}\right],$
where $Q\in^{n\times n}$ and $R\in^{m\times m}$ are the symmetric positive definite matrices that parameterize the cost. We require the pair $(A,B)$ to be stabilizable, and since $Q$ is positive definite (which we also show by $Q\succ 0$), the pair $(A,Q^{1/2})$ is observable, As a result, the unique optimal controller is a linear state-feedback $u^{\ast}_{t}=-K^{*}x_{t}$ where $K^{*}\in^{m\times n}$ is derived as follows
@@
and $P^{*}\in^{n\times n}$ denotes the unique positive definite solution to the undiscounted discrete-time algebraic Riccati equation (ARE) [2]:
@@
Moreover, following the notation in [22], we denote the $P^{*}$-induced norm by $\|X\|_{*}$. Since the optimal controller is a linear state feedback, we formulate the costs directly in terms of the feedback gain $K$, taking $u_{t}=-Kx_{t}$ for all $t\geq 0$. Throughout the paper, we use the terms “policy” and “feedback gain” interchangeably to refer to such matrices $K$.
We now introduce important notation used in the algorithm description and the proof of the main result. Let $N$ be the horizon length and $h$ the initial time step. The true finite-horizon cost $J_{h}(K_{h})$ of a policy $K_{h}$ is defined as
(5) $J_{h}(K_{h}):={\mathbb{E}}_{x_{0}\sim\mathcal{D}}\Bigg[\sum_{t=h+1}^{N-1}x_{t-h}^{\top}\left(Q+(K^{*}_{t})^{\top}RK^{*}_{t}\right)x_{t-h}+x_{0}^{\top}\left(Q+K_{h}^{\top}RK_{h}\right)x_{0}+x_{N-h}^{\top}Q_{N}x_{N-h}\Bigg],$
@@
  • $x_{0}$ denotes the initial state $x_{0}$ drawn from distribution $\mathcal{D}$,
  • $Q_{N}$ is the terminal cost matrix, which can be chosen arbitrarily,
@@
(6) $\widetilde{J}_{h}(K_{h}):={\mathbb{E}}_{x_{0}\sim\mathcal{D}}\Bigg[\sum_{t=h+1}^{N-1}x_{t-h}^{\top}\left(Q+\widetilde{K}_{t}^{\top}R\widetilde{K}_{t}\right)x_{t-h}+x_{0}^{\top}\left(Q+K_{h}^{\top}RK_{h}\right)x_{0}+x_{N-h}^{\top}Q_{N}x_{N-h}\Bigg],$
@@
$x_{1}=(A-BK_{h})x_{0}.$
The key difference between $\widetilde{J}_{h}(K_{h})$ and $J_{h}(K_{h})$ lies in the use of $\widetilde{K}_{t}$ versus $K^{*}_{t}$ for $t\in\{h+1,\dotsc,N-1\}$. This distinction implies that $\widetilde{J}_{h}(K_{h})$ incorporates all errors from earlier steps, precisely the ones at $\{N-1,\dotsc,h+1\}$, as the procedure is backward.
@@
(7) $\widetilde{J}_{h}(K_{h};x_{0})$ $:=\sum_{t=h+1}^{N-1}x_{t-h}^{\top}\left(Q+\widetilde{K}_{t}^{\top}R\widetilde{K}_{t}\right)x_{t-h}+x_{0}^{\top}(Q+K_{h}^{\top}RK_{h})x_{0}+x_{N-h}^{\top}Q_{N}x_{N-h}$
(8) $=x_{0}^{\top}(Q+K_{h}^{\top}RK_{h})x_{0}+x_{0}^{\top}(A-BK_{h})^{\top}\widetilde{P}_{h+1}(A-BK_{h})x_{0},\qquad$
@@
$\widetilde{J}_{h}(K_{h})={\mathbb{E}}_{x_{0}\sim\mathcal{D}}\left[\widetilde{J}_{h}(K_{h};x_{0})\right].$
Using (8), we can compute the gradient of $\widetilde{J}_{h}(K_{h};x_{0})$ with respect to $K_{h}$ as follows:
(9) $\nabla\widetilde{J}_{h}(K_{h};x_{0})$ $=\nabla\Bigg(x_{0}^{\top}K_{h}^{\top}RK_{h}x_{0}+x_{0}^{\top}K_{h}^{\top}B^{\top}\widetilde{P}_{h+1}BK_{h}x_{0}-2x_{0}^{\top}A^{\top}\widetilde{P}_{h+1}BK_{h}x_{0}\Bigg)$
(10) $=2RK_{h}x_{0}x_{0}^{\top}+2B^{\top}\widetilde{P}_{h+1}BK_{h}x_{0}x_{0}^{\top}-2B^{\top}\widetilde{P}_{h+1}Ax_{0}x_{0}^{\top}$
(11) $=2\left((R+B^{\top}\widetilde{P}_{h+1}B)K_{h}-B^{\top}\widetilde{P}_{h+1}A\right)x_{0}x_{0}^{\top},$
@@
(12) $\nabla\widetilde{J}_{h}(K_{h})$ $={\mathbb{E}}_{x_{0}\sim\mathcal{D}}\left[\nabla\widetilde{J}_{h}(K_{h};x_{0})\right]$
(13) $=2\left((R+B^{\top}\widetilde{P}_{h+1}B)K_{h}-B^{\top}\widetilde{P}_{h+1}A\right){\mathbb{E}}_{x_{0}\sim\mathcal{D}}\left[x_{0}x_{0}^{\top}\right]$
@@
$Q_{h}(x_{0},u_{0})$ $:=x_{0}^{\top}Qx_{0}+u_{0}^{\top}Ru_{0}+\sum_{t=h+1}^{N-1}x_{t-h}^{\top}\left(Q+\widetilde{K}_{t}^{\top}R\widetilde{K}_{t}\right)x_{t-h}+x_{N-h}^{\top}Q_{N}x_{N-h}$
$=x_{0}^{\top}Qx_{0}+u_{0}^{\top}Ru_{0}+\widetilde{J}_{h+1}(\widetilde{K}_{h+1};Ax_{0}+Bu_{0})$
(15) $=x_{0}^{\top}Qx_{0}+u_{0}^{\top}Ru_{0}+(Ax_{0}+Bu_{0})^{\top}\widetilde{P}_{h+1}(Ax_{0}+Bu_{0}),$
@@
$\widetilde{J}_{h}(K_{h})={\mathbb{E}}_{x_{0}\sim\mathcal{D}}\left[Q_{h}(x_{0},-K_{h}x_{0})\right].$
@@
(22) ${\mathbb{E}}_{x_{0}\sim{\mathcal{D}}}\left[x_{0}^{\top}\widetilde{P}_{h}x_{0}\right]$ $=\widetilde{J}_{h}(\widetilde{K}_{h}),$
(23) ${\mathbb{E}}_{x_{0}\sim{\mathcal{D}}}\left[x_{0}^{\top}\widetilde{P}^{*}_{h}x_{0}\right]$ $=\widetilde{J}_{h}(\widetilde{K}^{*}_{h}),\quad\text{and}$
(24) ${\mathbb{E}}_{x_{0}\sim{\mathcal{D}}}\left[x_{0}^{\top}P^{*}_{h}x_{0}\right]$ $=J_{h}(K^{*}_{h}).$
@@
We now introduce the following mild assumption:
@@
AA in (1) is non-singular.
Under this assumption, we have the following result from [19] which ensures that the Riccati operator $\mathcal{R}$ is non-expansive under $\delta(\cdot,\cdot)$: